OpenStudy (anonymous):

the first angle of a triangle is double the second angle.the third angle of the triangle is 12 degrees is less than the second angle.let the second angle=x

OpenStudy (anonymous):

first angle = 2x third angle = x-12 so x + 3x + (x - 12) = 180 - solve

OpenStudy (anonymous):

make shore u do the problime in the right order thow

OpenStudy (anonymous):

sorry thats x + 2x + x -12 = 180 4x = 192 x = 48 degrees = second angle i expect you can now find the other angles